# From Prompts to Protection: Large Language Model-Enabled In-Context Learning for Smart Public Safety UAV

### v1

@@ Abstract
A public safety Unmanned Aerial Vehicle (UAV) enhances situational awareness in emergency response. Its agility and ability to optimize mobility and establish Line-of-Sight (LoS) communication make it increasingly vital for managing emergencies such as disaster response, search and rescue, and wildfire monitoring. While Deep Reinforcement Learning (DRL) has been applied to optimize UAV navigation and control, its high training complexity, low sample efficiency, and simulation-to-reality gap limit its practicality in public safety. Recent advances in Large Language Models (LLMs) offer a compelling alternative. With strong reasoning and generalization capabilities, LLMs can adapt to new tasks through In-Context Learning (ICL), which enables task adaptation via natural language prompts and example-based guidance, without retraining. Deploying LLMs at the network edge, rather than in the cloud, further reduces latency and preserves data privacy, thereby making them suitable for real-time, mission-critical public safety UAVs. This paper proposes the integration of LLM-enabled ICL with public safety UAV to address the key functions, such as path planning and velocity control, in the context of emergency response. We present a case study on data collection scheduling where the LLM-enabled ICL framework can significantly reduce packet loss compared to conventional approaches, while also mitigating potential jailbreaking vulnerabilities. Finally, we discuss LLM optimizers and specify future research directions. The ICL framework enables adaptive, context-aware decision-making for public safety UAV, thus offering a lightweight and efficient solution for enhancing UAV autonomy and responsiveness in emergencies.
@@
Unmanned Aerial Vehicle, Large Language Models, In-Context Learning, Network Edge
@@ I Introduction
Unmanned Aerial Vehicles (UAVs) play a crucial role in various sectors, including energy, environmental monitoring, parcel delivery, and public safety. In particular, they are transformative in public safety applications, including post-disaster assessment, search and rescue missions, wildfire monitoring, border surveillance, and crime scene reconstruction, due to their agility, optimized mobility, and ability to establish reliable Line-of-Sight (LoS) communications. A public safety UAV enhances situational awareness for police, firefighters, and other first responders, thereby improving incident management and overall operational efficiency[1],[2].
@@
UAVs can leverage Machine Learning (ML) techniques for optimizing navigation and control, particularly in emergencies. For example, they employ Deep Reinforcement Learning (DRL) to optimize flight trajectories based on victims’ needs or to identify casualties and collect environmental data efficiently[3]. However, DRL solutions involve complex training processes and face challenges such as the simulation-to-reality gap and low sample efficiency, which are ill-suited to the urgent demands of emergency scenarios. Therefore, lightweight and adaptive solutions are required to quickly adapt to the emergency and enhance the success rate of rescue operations.
@@
Recent advancements in artificial intelligence, particularly in Large Language Models (LLMs), are driving significant transformations across UAV communications[4]. With strong reasoning and generalization capabilities, LLMs enable advanced understanding, flexible adaptation, and real-time responsiveness in diverse UAV scenarios. Integrating LLMs into UAV communications offers a promising path toward greater autonomy, equipping UAVs with enhanced decision-making abilities and enabling more effective responses during emergencies.
@@
The conventional approach to deploying LLMs relies on cloud-based infrastructures, which often introduces communication delays and raises privacy concerns, particularly problematic for time-sensitive and mission-critical UAV operations[5]. A recent trend is to shift LLM deployment to the network edge through edge server integration. This approach allows LLMs to directly serve the computational and operational needs of edge devices such as UAVs. Notably, LLMs can be taught new tasks using In-Context Learning (ICL) by the UAV, enabling rapid adaptation without retraining. For instance, ICL can facilitate key UAV functions, including path planning and velocity control, data collection scheduling, and power management, enhancing real-time decision-making in emergencies.
@@
Figure 1: Illustration of LLM-enabled ICL for a public safety UAV. The public safety UAV collects sensory data and submits it to the edge LLM, where a task description, which is a basis for ICL, is formulated with the following components: task goal, input data, rules, output, and feedback. The system prompt defines the rule of LLM, e.g., data collector.
@@
This paper proposes the use of LLM-enabled ICL to facilitate the key functions in public safety UAV. We systematically categorize these functions and present detailed task descriptions for each, demonstrating how ICL can enable the public safety UAV to adapt rapidly to emergency scenarios, improve decision-making, and enhance operational efficiency without requiring model retraining. As depicted in Fig. 1, ICL leverages structured natural language prompts, i.e., task descriptions, to encapsulate task objectives, while curated demonstration sets convey relevant system states and environmental conditions such as sensory data. This approach enables LLMs to infer an effective output for the task, without requiring explicit model retraining. By using contextually rich examples, ICL supports adaptive, on-the-fly decision-making across a broad range of UAV-Assisted Sensor Network (UASNETs) tasks, making it particularly valuable for time-sensitive applications such as emergency response.
The contributions of the paper are listed as follows:
  • A classification of key functions in public safety UAVs, such as path planning and velocity control. These functions can be effectively addressed using LLM-enabled ICL, providing a structured foundation for its application in emergencies. This classification serves as a foundational reference to guide future research and development in integrating LLM-enabled ICL into public safety UAVs, especially where DRL optimizers fall short.
  • A case study on data collection schedule, showcasing how an LLM-enabled ICL framework significantly reduces packet loss compared to conventional strategies, while also addressing potential jailbreaking vulnerabilities.
  • A discussion on LLM optimizers along with future research directions. One such direction involves the use of Diffusion Models (DMs) to generate diverse, high-quality synthetic examples that can be used for enriching task descriptions.
@@
The remainder of this paper is organized as follows: Section II discusses the deployment of LLMs. Section III provides background on LLM-enabled ICL and its effectiveness. Section IV explores the application of LLM-enabled ICL for public safety UAV and its various functions. Section V presents a case study on data collection schedule, along with investigating jailbreaking attacks. Section VI discusses LLM optimizers and their merits for public safety UAV. Section VII outlines future directions, including the use of DMs to enrich task descriptions. Finally, Section VIII concludes the paper.
Table I: Overview of In-Context Learning for LLMs in Public Safety UAV Applications.
Aspect Description
Definition ICL enables LLMs to learn from contextual information embedded in prompts containing task-specific instructions and demonstrations written in natural language.
Capabilities Allows LLMs to generalize across tasks by leveraging pre-trained knowledge and adapting to new tasks through example-based prompts. Considered a core feature of modern LLMs.
Key Considerations The design of demonstrations, including example selection, formatting, and ordering, strongly influences performance and should be carefully crafted.
Application in Public Safety ICL offers a lightweight method to facilitate key functions such as path planning by reusing prior solutions as input prompts, supporting fast adaptation in emergencies.
@@ II Deployment of LLMs
Thanks to their strong generalization and reasoning capabilities, LLMs enable UAVs to interpret complex, dynamic environments and develop coherent, sequential control strategies for navigation, surveillance, and decision-making tasks. These capabilities make LLMs well-suited for real-time operations in mission-critical scenarios such as disaster response, infrastructure inspection, and surveillance. However, most existing LLM deployments rely heavily on cloud-based infrastructures, which impose critical limitations in terms of latency, network bandwidth, and data privacy, making them suboptimal or even unfeasible for delay-sensitive or communication-limited UAV applications. To address this problem, efficient model compression techniques such as quantization, pruning, and knowledge distillation are used to minimize model size and computational load without compromising performance. In parallel, fast decoding strategies such as speculative decoding, early termination, parallel decoding, and sparse attention reduce inference latency and memory requirements during autoregressive token generation.
@@
Once LLMs are efficiently deployed at the edge, they can be leveraged for context-aware inference through prompt engineering. Techniques such as CoT prompting, prompt-based planning, self-refinement prompting, and ICL enable LLMs to perform complex tasks by interpreting instructions and demonstrations embedded in prompts without the need for parameter updates. ICL is particularly effective in generalizing tasks because it uses previously trained knowledge and adapts to examples, with performance depending on the design of the prompts, including the selection, arrangement, and formatting of the demonstrations. In dynamic, resource-constrained scenarios such as UASNETs, ICL offers tangible benefits as previous solutions can serve as reusable inputs for new challenges, increasing the flexibility and responsiveness of UAVs. [7].
III LLM-enabled In-Context Learning
This section provides background on LLM-enabled ICL and discusses its effectiveness. ICL refers to the ability of a pretrained language model to perform a task by conditioning on a sequence of input-output demonstration examples, without requiring any updates to the model parameters. Given an input query x and a set of candidate answers $\mathcal{Y}=\{y_{1},\ldots,y_{m}\}$, the model selects the output with the highest likelihood based on a context $C$, which includes an optional task instruction I and k demonstration examples. This context can be expressed as either:
$\mathcal{C}=\{I,s(x_{1},y_{1}),\ldots,s(x_{k},y_{k})\}$
where $s$ denotes a natural language representation of the examples.
ICL differs from related learning paradigms in the following ways:
@@ III LLM-enabled In-Context Learning
The likelihood of a candidate answer $y_{j}$ is computed using a scoring function $f$ is given by:
$P(y_{j}\mid x)\triangleq f_{M}(y_{j},\mathcal{C},x),$
and the final prediction $\hat{y}$ is given by:
$\hat{y}=\arg\max_{y_{j}\in\mathcal{Y}}P(y_{j}\mid x).$
Min et al.[9] reveal that the effectiveness of ICL relies more on the structural presentation of input and label spaces than on the correctness of individual demonstration labels. Surprisingly, replacing ground truth labels with random ones in demonstrations has only a marginal effect on performance. Moreover, models can retain up to 95% of performance gains using either inputs or label sets alone—provided the formatting is appropriate. These insights suggest that formatting and representational cues play a more critical role than previously assumed, with implications for both model design and future research in prompt engineering and meta-training for ICL. Table I provides an overview of LLM-enabled ICL for public safety UAV. The key point is that the way demonstrations are designed, such as the choice of examples, their formatting, and sequence, has a significant impact on performance and must be meticulously crafted.
IV LLM-enabled In-Context Learning for Optimized Public Safety UAV
This section highlights the potential of LLM-enabled ICL in optimizing key functions in public safety to dynamically adapt based on real-time context and prior experiences, all without the need for retraining the model. We consider a post-disaster environment where an LLM is hosted at the network edge. The UAV can physically approach individual ground sensors. The short, LoS-dominant communication link between a UAV and a ground sensor offers a significant channel gain and supports high-speed data transmission. In this context, employing a UAV for data collection enhances network throughput and extends coverage beyond what terrestrial gateways can achieve[10].
@@ IV-A Path Planning
ML is essential for UAV placement and path planning due to the complex nature of designing UASNETs and the critical requirements in applications such as public safety. Determining the optimal horizontal and vertical placement of the UAV, as well as planning its trajectories to interact with ground or flying objects, requires DRL optimizers that can handle large-scale optimization problems and dynamic environmental factors. In a nutshell, DRL is indispensable for UAV placement and path planning and facilitates complex optimization tasks, supports real-time adaptation to changing conditions, and enhances capabilities for detecting and responding to noncompliant UAV in critical applications like public safety. Path planning involves the optimization of continuous variables. In this work, we simplify the problem by discretizing the action space and restricting the UAV’s motion to a finite number of directions (e.g., north, south, east, and west). This abstraction not only reduces the computational complexity but also makes the problem accessible for LLM-enabled ICL. Starting with a task description depicted in Fig. 2(b), which defines the task and provides in-context demonstrations, the LLM as optimizer generates candidate solutions. These solutions are then implemented in the environment, and their associated costs are computed. Feedback is derived from the latest sensory data, the implemented solution, and the obtained cost. This feedback, along with updated sensory data, is fed back into the LLM. The process repeats iteratively until satisfactory solutions are generated, enabling the LLM to refine its outputs dynamically based on real-world performance.
@@ IV-B Velocity Control
As depicted in Fig. 3, the UAV hovers over the disaster area where it adjusts its velocity based on the requirements of the ground sensors. The slow mobility of a UAV can lead to buffer overflows at the ground sensors, as newly arriving data is not immediately captured by the UAV. The fast mobility of a UAV can harm channel conditions and increase communication failures. Onboard velocity control of the UAV is essential to minimize packet loss due to ground sensor buffer overflows and communication failures[12]. An LLM-enabled ICL framework can be employed to optimize the velocity of the UAV to minimize the overall packet loss. Starting from a task description depicted in Fig. 2(a), which defines the velocity control task and provides in-context examples, the LLM acts as an optimizer to generate candidate velocities. The performance of the candidate velocities is evaluated by calculating a cost metric. The feedback is then shaped by the latest sensor data, velocity, together with the calculated cost. This feedback is fed back to the LLM along with the current sensor data. The LLM iteratively refines its control strategies using this closed-loop process until the system converges to stable, optimal velocity tracking.

### v2

# From Prompts to Protection: Large Language Model-assisted In-Context Learning for Smart Public Safety UAV

@@ Abstract
A public safety Uncrewed Aerial Vehicle (UAV) enhances situational awareness during emergency response. Its agility, mobility optimization, and ability to establish Line-of-Sight (LoS) communication make it increasingly important for managing emergencies such as disaster response, search and rescue, and wildfire monitoring. Although Deep Reinforcement Learning (DRL) has been used to optimize UAV navigation and control, its high training complexity, low sample efficiency, and the simulation-to-reality gap limit its practicality in public safety applications. Recent advances in Large Language Models (LLMs) present a promising alternative. With strong reasoning and generalization abilities, LLMs can adapt to new tasks through In-Context Learning (ICL), enabling task adaptation via natural language prompts and example-based guidance without retraining. Deploying LLMs at the network edge, rather than in the cloud, further reduces latency and preserves data privacy, making them suitable for real-time, mission-critical public safety UAVs. This paper proposes integrating LLM-assisted ICL with public safety UAVs to address key functions such as path planning and velocity control in emergency response. We present a case study on data collection scheduling, demonstrating that the LLM-assisted ICL framework can significantly reduce packet loss compared to conventional approaches while also mitigating potential jailbreaking vulnerabilities. Finally, we discuss LLM optimizers and outline future research directions. The ICL framework enables adaptive, context-aware decision-making for public safety UAVs, offering a lightweight and efficient solution to enhance UAV autonomy and responsiveness in emergencies.
@@
Uncrewed Aerial Vehicle, Large Language Models, In-Context Learning, Network Edge, Public Safety
@@ I Introduction
Uncrewed Aerial Vehicles (UAVs) play a crucial role in sectors such as energy, environmental monitoring, parcel delivery, and public safety. They are especially transformative in public safety applications, including post-disaster assessment, search and rescue missions, wildfire monitoring, border surveillance, and crime scene reconstruction, due to their agility, optimized mobility, and ability to establish reliable Line-of-Sight (LoS) communications. Public safety UAVs enhance situational awareness for police, firefighters, and other first responders, thereby improving incident management and overall operational efficiency.[11].
@@
UAVs can use Machine Learning (ML) techniques to optimize navigation and control, especially during emergencies. For example, they apply Deep Reinforcement Learning (DRL) to optimize flight trajectories based on victims’ needs or to identify casualties and efficiently collect environmental data[7]. However, DRL solutions require complex training and face challenges such as the simulation-to-reality gap and low sample efficiency, making them unsuitable for the urgent demands of emergency scenarios. Therefore, lightweight and adaptive solutions are needed to respond effectively to emergencies and improve the success rate of rescue operations.
@@
Recent advancements in artificial intelligence, particularly in Large Language Models (LLMs), are driving significant transformations in UAV communications [6]. With strong reasoning and generalization capabilities, LLMs enable advanced understanding, flexible adaptation, and real-time responsiveness in diverse UAV scenarios [12]. Integrating LLMs into UAV communications offers a promising path toward greater autonomy, equipping UAVs with enhanced decision-making abilities and enabling more effective responses during emergencies.
@@
The conventional approach to deploying LLMs relies on cloud-based infrastructures, which often introduces communication delays and raises privacy concerns – issues that are particularly problematic for time-sensitive and mission-critical UAV operations[13]. A recent trend is to shift LLM deployment to the network edge through edge server integration. This approach enables LLMs to directly serve the computational and operational needs of edge devices such as UAVs. Notably, LLMs can learn new tasks using In-Context Learning (ICL) by the UAV, allowing rapid adaptation without retraining. For example, ICL can support key UAV functions such as path planning, velocity control, data collection scheduling, and power management, thereby enhancing real-time decision-making in emergencies. [14].
ICL is inherently lightweight because it eliminates the need for conventional model retraining and complex optimization. Unlike DRL or supervised methods that require extensive gradient updates, ICL adapts to new environments purely through prompt design and contextual demonstrations, minimizing computational and energy costs. Its training-free nature allows real-time responsiveness to dynamic UAVs or network conditions without simulation cycles. Moreover, ICL is compatible with compact, instruction-tuned models that can be quantized and deployed efficiently on edge devices, enabling low-latency reasoning at the network edge. By leveraging natural-language task formulations instead of handcrafted reward functions, ICL offers a simplified, interpretable, and resource-efficient alternative to conventional learning pipelines.
@@
Figure 1: Illustration of LLM-assisted ICL for a public safety UAV. The public safety UAV collects sensory data and submits it to the edge LLM, where a task description, which is a basis for ICL, is formulated with the following components: task goal, input data, rules, output, and feedback. The system prompt defines the rule of LLM, e.g., data collector.
@@
This paper advances the state of LLM-assisted UAV intelligence from conceptual discussions and single-task prototypes toward a fully deployable, network-aware control framework. Unlike prior studies that focus on descriptive overviews [6], [15], [10] or isolated scheduling scenarios [2], the proposed approach integrates in-network ICL with verifier-guided optimization to achieve safe, adaptive, and scalable multiple UAV functions. By coupling telemetry-driven prompt construction with a unified safety-verifier, the framework enables real-time context adaptation under dynamic wireless conditions while maintaining energy efficiency and decision robustness. Through quantitative evaluation across multiple performance metrics, this paper bridges the gap between reasoning-level intelligence and network-level optimization, demonstrating the practical feasibility of LLM-assisted UAV autonomy in real communication environments.As depicted in Fig. 1, ICL leverages structured natural language prompts, i.e., task descriptions, to encapsulate task objectives, while curated demonstration sets convey relevant system states and environmental conditions such as sensory data. This approach enables LLMs to infer an effective output for the task, without requiring explicit model retraining. By using contextually rich examples, ICL supports adaptive, on-the-fly decision-making across a broad range of UAV-Assisted Sensor Network (UASNETs) tasks, making it particularly valuable for time-sensitive applications such as emergency response.
The contributions of this paper are summarized as follows:
Firstly, we propose a unified framework and functional taxonomy for LLM-assisted ICL in public-safety UAVs. Unlike prior studies focusing on individual tasks such as data collection scheduling, this work establishes an integrated architecture that jointly supports key UAV functions such as path planning, velocity control, and power management. The framework provides a structured foundation and standardized blueprint for applying ICL in emergency contexts, particularly where DRL-based optimizers face high training complexity and limited real-world transferability. Secondly, we present a case study on data collection scheduling that demonstrates the effectiveness and security of the proposed ICL framework. Through a black-box attacker model, we analyze context-based jailbreaking threats and show that the ICL-based Data Collection Scheduling (ICLDC) system dynamically adapts to network variations while maintaining low packet loss and stable scheduling performance. This analysis highlights both the operational benefits and the resilience of LLM-assisted decision-making in mission-critical UAV environments. Thirdly, we provide a comparative discussion of optimization paradigms for public-safety UAVs, covering DRL, Diffusion Model (DM)+DRL hybrids, and LLM-assisted ICL. We further outline key research trends, including DM-based synthetic data generation, on-device LLM deployment, collaborative reasoning for UAV swarms, and real-world validation across domains. Together, these insights offer a practical roadmap for deploying LLM-assisted optimization in safety-critical UAV networks, balancing adaptability, efficiency, and robustness.
@@
The remainder of this paper is organized as follows: Section II discusses the deployment of LLMs. Section III provides background on LLM-assisted ICL and its effectiveness. Section IV explores the application of LLM-assisted ICL for public safety UAV and its various functions. Section V presents a case study on data collection schedule, along with investigating jailbreaking attacks. Section VI discusses LLM optimizers and their merits for public safety UAV. Section VII outlines future directions, including the use of DMs to enrich task descriptions. Finally, Section VIII concludes the paper.
@@ II Deployment of LLMs
Thanks to their strong generalization and reasoning capabilities, LLMs enable UAVs to interpret complex, dynamic environments and develop coherent control strategies for navigation, surveillance, and decision-making tasks. These capabilities make LLMs well-suited for real-time operations in mission-critical scenarios such as disaster response, infrastructure inspection, and surveillance. However, most existing LLM deployments rely heavily on cloud-based infrastructures, which impose critical limitations in terms of latency, network bandwidth, and data privacy, making them suboptimal or even unfeasible for delay-sensitive or communication-limited UAV applications. To address this problem, efficient model compression techniques such as quantization, pruning, and knowledge distillation are used to minimize model size and computational load without compromising performance. In parallel, fast decoding strategies such as speculative decoding, early termination, parallel decoding, and sparse attention reduce inference latency and memory requirements.
@@
Once LLMs are efficiently deployed at the edge, they can be leveraged for context-aware inference through prompt engineering. Techniques such as CoT prompting, prompt-based planning, self-refinement prompting, and ICL enable LLMs to perform complex tasks by interpreting instructions and demonstrations embedded in prompts without the need for parameter updates. ICL is particularly effective in generalizing tasks because it uses previously trained knowledge and adapts to examples, with performance depending on the design of the prompts, including the selection, arrangement, and formatting of the demonstrations. In dynamic, resource-constrained scenarios such as UASNETs, ICL offers tangible benefits as previous solutions can serve as reusable inputs for new challenges, increasing the flexibility and responsiveness of UAVs[10].
III LLM-assisted In-Context Learning
This section provides background on LLM-assisted ICL and discusses its effectiveness. The context used for ICL consists of an optional task instruction and a sequence of demonstration examples. The task instruction specifies the overall goal or nature of the task for the model. Each demonstration example consists of an input and its corresponding output, represented in natural language so that the model can interpret it. By conditioning on this context, the pretrained language model can perform a new task without updating its parameters.
The examples in the context provide patterns for the model to follow, allowing it to generate an appropriate output for a new input query based on the information provided.
@@ III LLM-assisted In-Context Learning
The likelihood of each candidate answer is determined using a scoring function, which evaluates how well the candidate aligns with the input query and the context provided by the task instruction and demonstration examples. Based on these scores, the model selects the candidate answer with the highest likelihood as its final prediction. In other words, the model compares all possible outputs in light of the context and chooses the one that best fits the pattern established by the examples and instructions, enabling accurate task completion without updating model parameters.
Min et al.[9] reveal that the effectiveness of ICL relies more on the structural presentation of input and label spaces than on the correctness of individual demonstration labels. Surprisingly, replacing ground truth labels with random ones in demonstrations has only a marginal effect on performance. Moreover, models can retain up to 95% of performance gains using either inputs or label sets alone—provided the formatting is appropriate. These insights suggest that formatting and representational cues play a more critical role than previously assumed, with implications for both model design and future research in prompt engineering and meta-training for ICL.
IV LLM-assisted In-Context Learning for Optimized Public Safety UAV
This section highlights the potential of LLM-assisted ICL in optimizing key functions in public safety to dynamically adapt based on real-time context and prior experiences, all without the need for retraining the model, and provide discussion. We consider a post-disaster environment where an LLM is hosted at the network edge. The UAV can physically approach individual ground sensors. The short, LoS-dominant communication link between a UAV and a ground sensor offers a significant channel gain and supports high-speed data transmission. In this context, employing a UAV for data collection enhances network throughput and extends coverage beyond what terrestrial gateways can achieve[3].
Throughout this section, the term “feedback” refers to externally measured performance indicators, such as packet loss, throughput, and energy consumption, which are computed by the UAV or the edge controller after each LLM-generated decision. These measurements are then summarized as contextual information and appended to the next prompt for further reasoning. The LLM itself does not calculate the numerical cost; instead, all costs are derived from observable system telemetry without relying on any ground-truth or oracle data.
@@ IV-A Path Planning
ML is essential for UAV placement and path planning due to the complex nature of designing UASNETs and the critical requirements in applications such as public safety. Determining the optimal horizontal and vertical placement of the UAV, as well as planning its trajectories to interact with ground or flying objects, requires DRL optimizers that can handle large-scale optimization problems and dynamic environmental factors. In a nutshell, DRL is indispensable for UAV placement and path planning and facilitates complex optimization tasks, supports real-time adaptation to changing conditions, and enhances capabilities for detecting and responding to noncompliant UAV in critical applications like public safety. Path planning involves the optimization of continuous variables. In this work, we simplify the problem by discretizing the action space and restricting the UAV’s motion to a finite number of directions (e.g., north, south, east, and west). The use of a four-direction discrete action model prioritizes rapid, adaptive decision-making over aerodynamic precision, an essential trade-off in disaster-response missions. By constraining the action space to cardinal directions, the framework ensures computational tractability and real-time responsiveness for LLM-assisted reasoning, even on resource-limited UAV or edge platforms. This abstraction aligns with real-world UAV architectures, where high-level planners issue directional or waypoint commands while low-level controllers manage flight dynamics. Consequently, the discrete model serves as a practical and efficient bridge between natural-language reasoning and executable UAV control, enabling fast, mission-relevant adaptation in dynamic and safety-critical environments. Starting with a task description depicted in Fig. 2(b), which defines the task and provides in-context demonstrations, the LLM as an optimizer generates candidate solutions. These solutions are then implemented in the environment, and their associated costs are computed. Feedback is derived from the latest sensory data, the implemented solution, and the obtained cost. This feedback, along with updated sensory data, is fed back into the LLM. The process repeats iteratively until satisfactory solutions are generated, enabling the LLM to refine its outputs dynamically based on real-world performance.
In the path-planning task, the cost is defined as a combination of observable performance indicators, including the travel distance of the UAV, the measured packet-loss rate on the communication links, and the number of sensors whose data queues exceed a defined threshold. These values are obtained directly from UAV odometry, link-level acknowledgements, and queue reports from ground sensors, and then summarized in natural language as part of the feedback context for subsequent reasoning.
@@ IV-B Velocity Control
As depicted in Fig. 3, the UAV hovers over the disaster area where it adjusts its velocity based on the requirements of the ground sensors. The slow mobility of a UAV can lead to buffer overflows at the ground sensors, as newly arriving data is not immediately captured by the UAV. The fast mobility of a UAV can harm channel conditions and increase communication failures. Onboard velocity control of the UAV is essential to minimize packet loss due to ground sensor buffer overflows and communication failures[5]. An LLM-assisted ICL framework can be employed to optimize the velocity of the UAV to minimize the overall packet loss. Starting from a task description depicted in Fig. 2(a), which defines the velocity control task and provides in-context examples, the LLM acts as an optimizer to generate candidate velocities. The performance of the candidate velocities is evaluated by calculating a cost metric. The feedback is then shaped by the latest sensor data, the current velocity, and the calculated cost. This feedback is fed back to the LLM along with the current sensor data. The LLM iteratively refines its control strategies using this closed-loop process until the system converges to stable, optimal velocity tracking. For velocity control, the cost metric emphasizes queue stability and communication reliability. It is derived from the observed frequency of queue overflows, the rate of transmission failures recorded through link acknowledgments, and a small penalty associated with abrupt changes in velocity. These values are collected by the UAV-edge controller and reported back as structured feedback, allowing the LLM to progressively refine its velocity adjustment policy.